# On distinguishing coloring and some variants of proper coloring of graphs derived from subdivision operations

### v1

@@ 3. Distinguishing number and distinguishing index
    (Kalinowski, Pilśniak, and Woźniak; [15, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
@@
    (Alikhani and Soltani; [4, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
@@
    (Kalinowski and Pilśniak; [16, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
@@
    (Alikhani and Soltani; [3, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
@@
    (Kalinowski and Pilśniak; [16]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is $C_{3},C_{4},$ or $C_{5}$.
@@
Since $G$ is connected, we have $\Delta(G)\geq 2$. Hence, we have at least $\lceil\sqrt{2}\rceil=2$ colors available to color $C(G)$. For $k\geq 0$, let $S_{k}(x)$ denote a sphere of radius $k$ with a center $x$ in $G$, that is, the set of all vertices at distance $k$ from $x$.
@@
If $n\geq 6$, then $D^{\prime}(G)=2\leq\lceil\sqrt{\Delta(G)}\rceil$ by Fact 3.1(3). We assume $n$ is either $4$ or $5$. In both cases, we can color the edges of $C(G)$ with two available colors so that the edge coloring is preserved only by $id_{G}$ (see Fig. 2).
$C(C_{4})$$C(C_{5})$$C(K_{4})$$C(K_{5})$
Figure 2. Edge colorings of $G$ with $2$ colors that are only preserved by $id_{G}$ where $G\in\{C(C_{4}),C(C_{5}),C(K_{4}),C(K_{5})\}$.
@@
$X_{v,u}=(\{v,w_{v,u}\},\{w_{v,u},u\})$ if $w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
@@
$N(v)=\{X_{v,u}:w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
@@
  • We pick $v_{1},v_{2}\in S_{1}(v_{0})$ and assign the pair $(1,1)$ of colors to $X_{v_{0},v_{1}}$ and $X_{v_{0},v_{2}}$. Since the degree of $v_{0}$ is at most $\Delta(G)$ and the pair $(1,1)$ is used twice to color the elements of $N(v_{0})$, we can color every pair of $N(v_{0})\setminus\{X_{v_{0},v_{1}},X_{v_{0},v_{2}}\}$ with different pairs of colors from $(1,1)$. We modify an algorithm due to Kalinowski–Pilśniak–Woźniak [15] to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$. Hence $v_{0}$ will be fixed by every color-preserving automorphism, and all vertices in $N(v_{0})$ will be fixed, except possibly $v_{1}$ and $v_{2}$. For $i\in\{1,2\}$, let
@@
[…]
      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$. Thus, $\{v_{0},u\}\in E(\overline{G})\subseteq E(C(G))$. We color all such edges with 1. On the other hand, we color all edges of the set $\{\{v_{0},u\}\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$ with 00.
@@
[…]
@@
Step 2: We color all the remaining edges in $C(G)$ with 00.
Step 3: Using Lemmas 2.3 and 2.4, we prove that the coloring that we defined above (say $c$) is preserved only by $id_{C(G)}$.
  • Let $\psi\in Aut(C(G))$ so that $\psi$ preserves $c$. By Lemma 2.4, $\psi(V_{i})=V_{i}$ for $i\in\{1,2\}$. Since all vertices of $T$ are fixed by $\psi$, we have $\psi(u)=u$ for all $u\in V_{1}$. By Lemma 2.3, $\psi$ is $id_{C(G)}$.
@@
  1. Case A: If $G$ has a central vertex $v_{0}$, then every element of $N(v_{0})$ obtains a different pair of colors (where $N(v)$ is defined in $Case(ii)$ for any vertex $v$ in the tree $G$). This is possible as we have at least $\Delta(G)$ different pairs of colors available. Hence, all vertices adjacent to $v_{0}$ in $G$ are fixed by every automorphism of $G$, preserving this coloring. If $v\in S_{k}(v_{0})$, then all the elements of $N(v)\backslash\{X_{v,u}:u\in S_{k-1}(v_{0})\}$ obtain a different pair of colors. By induction on the distance from $v_{0}$, all vertices of $G$ are fixed.
  2. Case B: If $G$ has a central edge $e_{0}=\{x,y\}$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$. If we put different colors on the edges $\{x,w_{x,y}\}$ and $\{w_{x,y},y\}$, then $x$ and $y$ are fixed by every automorphism. We color $T_{i}$ using the same method as in Case A for $i\in\{1,2\}$.
@@
In view of Fact 3.1(1), it suffices to show that $D(C(G))\leq D^{\prime\prime}(G)$. Let $f:V(G)\cup E(G)\to\{1,...,D^{\prime\prime}(G)\}$ be a total coloring of $G$ (not necessarily proper) that is preserved only by $id_{G}$. We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{x,y})=f(\{x,y\})$ for all $w_{x,y}\in V_{2}$. We claim that $\tilde{f}$ is preserved only by $id_{C(G)}$. Let $\psi$ be a $\tilde{f}$-preserving automorphism of $C(G)$. If $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$ where $\pi_{\alpha}$ is the induced automorphism of $\alpha\in Aut(G)$, which is defined in the proof of Lemma 2.5, then $H=Aut(C(G))\cong Aut(G)$ (see the proof of Lemma 2.5). Thus, $\psi=\pi_{\alpha}$ for some $\alpha\in Aut(G)$ where $\psi(V_{1})=V_{1}$ and $\psi(V_{2})=V_{2}$. Since $\psi$ is a $\tilde{f}$-preserving automorphism of $C(G)$, we can see that $\alpha$ preserves $f$ for the following reasons:
@@
    if $v\in V_{1}$, then $\tilde{f}(\psi(v))=\tilde{f}(v)\implies f(\psi(v))=f(v)\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v)$.
@@
    if $w_{x,y}\in V_{2}$, then $\tilde{f}(\psi(w_{x,y}))=\tilde{f}(w_{x,y})\implies\tilde{f}(\pi_{\alpha}(w_{x,y}))=f(\{x,y\})\implies\tilde{f}(w_{\alpha(x),\alpha(y)})=f(\{x,y\})\implies f(\{\alpha(x),\alpha(y)\})=f(\{x,y\})\implies f(\alpha(\{x,y\}))=f(\{x,y\})$.
@@
For any integer $n\geq 1$, we have $D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil$. This shows that the bounds mentioned in Theorems 3.2 and 3.4 are sharp (see Fig. 3).
@@
(1). By Facts 3.1(4) and 2.7 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
Case(i): $G\not\cong C_{n}$ for any $n\geq 3$. We show that $D^{\prime}(G^{+})\leq D^{\prime}(G)$. Then $D^{\prime}(G^{+})\leq\Delta(G)$ by Fact 3.1(5). Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$ and $E$ be the set of all endline edges of $G^{+}$. Define the mapping $g^{\prime}:E(G^{+})\rightarrow\{1,...,D^{\prime}(G)\}$ as follows:
@@
Case(ii): $G\cong C_{n}$ for some $n\geq 3$. Then $D(M(G))=D^{\prime}(G^{+})=2=\Delta(G)$. This also shows that the bound mentioned in (1) is sharp.
(2). Since every line graph is claw-free, and the distinguishing index of a connected claw-free graph is at most $3$ (see [29, Theorem 3.4]), we have $D^{\prime}(M(G))=D^{\prime}(L(G^{+}))\leq 3$. ∎
4. Total distinguishing chromatic number
Definition 4.1.
A Latin square of order $k$ is a $k\times k$ array based on the elements $1,2,...,k$ such that each element occurs exactly once in each row and exactly once in each column. A Latin square $M=[m_{i,j}]$ of order $k$ is said to be commutative if $m_{i,j}=m_{j,i}$, for $1\leq i,j\leq k$ and $M$ is said to be idempotent if $m_{i,i}=i$, for $1\leq i\leq k$. If the rows of $M$ are just cyclic permutations (one shift of the elements to the left) of the previous row, then $M$ is said to be anti-circulant.
1 5 2 6 3 7 4
5 2 6 3 7 4 1
2 6 3 7 4 1 5
6 3 7 4 1 5 2
3 7 4 1 5 2 6
7 4 1 5 2 6 3
4 1 5 2 6 3 7
Table 1. An anti-circulant idempotent commutative Latin square of order $7$.
Definition 4.2.
A proper edge coloring of a graph $G=(V(G),E(G))$ is a coloring of $E(G)$ so that any two adjacent edges have different colors. The chromatic index of $G$, denoted by $\chi^{\prime}(G)$, is the least number of colors needed for a proper edge coloring of $G$. Fix an integer $k$. We say that $G$ is $k$-edge choosable if for any assignment $L={L(e)}_{e\in E(G)}$ of the lists of available colors to the edges of $G$, there is a proper edge coloring $f$ of $G$ so that $f(e)\in L(e)$ and $|L(e)|=k$ for all $e\in E(G)$. The list chromatic index of $G$, denoted by $\chi_{L}^{\prime}(G)$, is the minimum integer $k$ such that $G$ is $k$-edge choosable.
Fact 4.3.
The following holds:
  1. (1)
    (Galvin; [14]) If $G$ is a bipartite graph, then $\chi^{\prime}(G)=\chi_{L}^{\prime}(G)=\Delta(G)$.
  2. (2)
    An idempotent commutative Latin square (ICLS) of order $k$ exists if and only if $k$ is odd.
  3. (3)
    If $M(k)=[m_{i,j}]$ is a Latin square where $m_{i,j}\equiv(i+j)k$ (mod $12k-$ 1), $11\leq m_{i,j}\leq 2k-$ 1, for $11\leq i,j\leq 2k-$ 1, then $M(k)$ is an anti-circulant ICLS of order $12k-$ 1.
  4. (4)
    (Xie–He;[35, Theorem 1]) If $G$ is a regular graph of even order $n$ and $\Delta(G)\geq\frac{2n}{3}+\frac{23}{6}$, then $G$ has TCC.
  5. (5)
    (Kalinowski–Pilśniak–Woźniak;[15]) Every connected graph $G$ satisfies the inequality $\chi^{\prime\prime}_{D}(G)\leq\chi^{\prime\prime}(G)+1$. Furthermore, $\chi_{D}^{\prime\prime}(G)=\chi^{\prime\prime}(G)$ if $\chi^{\prime\prime}(G)\geq\Delta(G)+2$.
4.1. Central graphs
Proposition 4.4.
If $G$ is a connected graph, then $\chi^{\prime\prime}_{D}(C(G))$ is $\Delta(C(G))+1$ or $\Delta(C(G))+2$.

### v2

@@ 3. Distinguishing number and distinguishing index
    (Kalinowski, Pilśniak, and Woźniak; [9, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
@@
    (Alikhani and Soltani; [3, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
@@
    (Kalinowski and Pilśniak; [10, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
@@
    (Alikhani and Soltani; [2, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
@@
    (Kalinowski and Pilśniak; [10]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is isomorphic to $C_{3},C_{4},$ or $C_{5}$.
Proposition 3.2.
For any integer $n\geq 4$, we have $D^{\prime}(C(K_{n}))=2$.
Proof.
By Lemma 2.3, we have $Aut(C(K_{n}))\cong Aut(K_{n})\cong S_{n}$. Thus, $|Aut(C(K_{n}))|=n!>1$ and so $C(K_{n})$ is not asymmetric. Consequently, $D^{\prime}(C(K_{n}))>1$ and it suffices to construct an edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that the only automorphism of $C(K_{n})$ that preserves $c$ is $id_{C(K_{n})}$. Let $V(K_{n})=\{v_{1},\dots,v_{n}\}$. Recall that for each edge $v_{i}v_{j}$ of $K_{n}$, $w_{v_{i}v_{j}}$ is the subdivision vertex introduced on that edge. Define the edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that for each $v_{i}\in V_{1}$, the incident edges $v_{i}w_{v_{i}v_{j}}$ are colored $2$ if $j<i$ and $1$ if $j>i$ (see Figure 2).
[…]
Figure 2. A schematic representation of the edge coloring of $C(K_{n})$. Edges $v_{i}w_{v_{i}v_{j}}$ are colored 22 (blue) if $j<i$ and 11 (red) if $j>i$.
Let $\varphi\in\mathrm{Aut}(C(K_{n}))$ be an automorphism that preserves $c$. For each $i\in\{1,...,n\}$, let $d_{2}(v_{i})$ denote the number of edges incident to $v_{i}$ that are colored with $2$. By construction, $d_{2}(v_{i})=i-1\text{ for all }i\in\{1,\dots,n\}.$ Hence, the values $d_{2}(v_{i})$ are distinct. Since $\varphi$ preserves $c$, we have $\varphi(v_{i})=v_{i}$ for all $v_{i}\in V(K_{n})$. Thus, by Lemma 2.1, $\varphi=id_{C(K_{n})}$. ∎
Proposition 3.3.
For any integer $n\geq 4$, we have $D^{\prime}(C(C_{n}))=2$.
Proof.
Since $Aut(C(C_{n}))\cong Aut(C_{n})\cong D_{n}$ by Lemma 2.3 where $D_{n}$ is the Dihedral group, the graph $C(C_{n})$ is not asymmetric, and thus $D^{\prime}(C(C_{n}))>1$. We show $D^{\prime}(C(C_{n}))\leq 2$. Let $V(C(G))=V_{1}\cup V_{2}$, where $V_{1}=V(C_{n})$ and $V_{2}=\{w_{xy}:xy\in E(C_{n})\}$ be the set of subdivision vertices. Since $n\geq 4$, we can select distinct $v_{1},v_{2},v_{3}\in V_{1}$ so that $v_{1}v_{2},v_{2}v_{3}\in E(C_{n})$. Define an edge coloring $c:E(C(C_{n}))\to\{0,1\}$ as follows (see Figure 3):
  1. (1)
    $c(v_{1}w_{v_{1}v_{2}})=1$ and $c(w_{v_{1}v_{2}}v_{2})=1$.
  2. (2)
    $c(v_{3}w_{v_{2}v_{3}})=1$.
  3. (3)
    $c(e)=0$ otherwise.
Let $\phi\in Aut(C(C_{n}))$ be an automorphism that preserves $c$. Let $d_{1}(u)$ denote the number of edges incident to $u$ colored with $1$. We can observe the following:
  • The vertices $w_{v_{1}v_{2}}$ and $w_{v_{2}v_{3}}$ are the unique vertices with $d_{1}(w_{v_{1}v_{2}})=2$ and $d_{1}(w_{v_{2}v_{3}})=1$. So $\phi(w_{v_{1}v_{2}})=w_{v_{1}v_{2}}$ and $\phi(w_{v_{2}v_{3}})=w_{v_{2}v_{3}}$. Since automorphism preserves neighbors, we have $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$ and $\phi(\{v_{2},v_{3}\})=\{v_{2},v_{3}\}$.
  • Among the neighbors $v_{2}$ and $v_{3}$ of $w_{v_{2}v_{3}}$, $v_{3}$ is the only vertex with $d_{1}(v_{3})=1$.111note that $d_{1}(v_{2})$ also is $1$, but $v_{2}\in N(w_{v_{1}v_{2}})$ Thus $\phi(v_{3})=v_{3}$. Consequently, $\phi(v_{2})=v_{2}$. Since $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$, we have $\phi(v_{1})=v_{1}$.
Since $\phi$ fixes three consecutive vertices $v_{1},v_{2},$ and $v_{3}$ of the cycle $C_{n}$, $\phi$ must fix all vertices of $C_{n}$. By Lemma 2.1, $\phi=id_{C(C_{n})}$. ∎
[…]
Figure 3. A schematic representation of the edge coloring of $C(C_{n})$. The edges $v_{1}w_{v_{1}v_{2}},w_{v_{1}v_{2}}v_{2}$, and $v_{3}w_{v_{2}v_{3}}$ are colored with 11 (red), and the rest of the edges are colored with 0 (black).
@@
Since $G$ is connected, we have $\Delta(G)\geq 2$. Hence,
$\left\lceil\sqrt{\Delta(G)}\right\rceil\geq\left\lceil\sqrt{2}\right\rceil=2.$
For $k\geq 0$, let $S_{k}(x)$ denote a sphere of radius $k$ with a center $x$ in $G$, that is, the set of all vertices at distance $k$ from $x$.
@@
By Propositions 3.2 and 3.3, $D^{\prime}(C(G))=2$. Since $\left\lceil\sqrt{\Delta(G)}\right\rceil\geq 2,$ we obtain $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil$.
@@
$X_{v,u}=(\{v,w_{vu}\},\{w_{vu},u\})$ if $w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
@@
$N(v)=\{X_{v,u}:w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
@@
  • We pick $v_{1},v_{2}\in S_{1}(v_{0})$ and assign the pair $(1,1)$ of colors to $X_{v_{0},v_{1}}$ and $X_{v_{0},v_{2}}$. Without loss of generality, we choose $v_{1}\in S_{1}(v_{0})$ such that $v_{1}$ has a neighbor in $S_{2}(v_{0})$. Since the degree of $v_{0}$ is at most $\Delta(G)$ and the pair $(1,1)$ is used twice to color the elements of $N(v_{0})$, we can color every pair of $N(v_{0})\setminus\{X_{v_{0},v_{1}},X_{v_{0},v_{2}}\}$ with different pairs of colors from $(1,1)$. Inspired by an algorithm due to Kalinowski–Pilśniak–Woźniak [9, Proof of Theorem 2.2], we modify the algorithm to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$. Hence $v_{0}$ will be fixed by every color-preserving automorphism, and all vertices in $N(v_{0})$ will be fixed, except possibly $v_{1}$ and $v_{2}$. For $i\in\{1,2\}$, let
@@
    Note that $T_{v_{1}}\neq\emptyset$ since $v_{1}$ has a neighbor in $S_{2}(v_{0})$, while $T_{v_{2}}$ may possibly be empty.
@@
    • ($\ast\ast$)
      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$. Hence $v_{0}u\in E(\overline{G})\subseteq E(C(G))$. Let
      $Z_{v_{1}}=\{v_{0}u\in E(C(G)):X_{v_{1},u}\in T_{v_{1}}\}.$
      We color all edges of $Z_{v_{1}}$ with $1$.
      Similarly, we color all edges of the set
      $Z_{v_{2}}=\{v_{0}u\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$
      with 0 (if this set is nonempty). In both cases (i.e., $Z_{v_{2}}=\emptyset$ or $Z_{v_{2}}\neq\emptyset$), the resulting color pattern in $C(G)$ distinguishes $v_{1}$ and $v_{2}$.
@@
Otherwise, $\phi(v_{1})=v_{2}$. Therefore, if $u\in S_{2}(v_{0})\cap N_{T}(v_{1})$, where $N_{T}(v)$ is the open neighborhood of a vertex $v$ in the tree $T$, then $\phi(u)\in S_{2}(v_{0})\cap N_{T}(v_{2})$. Since $uv_{0}\in E(C(G))$ and $f(uv_{0})=1$, we have $\phi(u)\phi(v_{0})\in E(C(G))$ and $f(\phi(u)\phi(v_{0}))=f(\phi(u)v_{0})=1$. This contradicts $(\ast\ast)$. ∎
@@
Step 2: We color all the remaining edges in $C(G)$ with 0.
Step 3: Using Lemmas 2.1 and 2.2, we prove that the coloring we defined above (say, $c$) is preserved only by $id_{C(G)}$.
  • Let $\psi\in Aut(C(G))$ be an automorphism that preserves $c$. By Lemma 2.2, $\psi(V_{i})=V_{i}$ for $i\in\{1,2\}$. Since all vertices of $T$ are fixed by $\psi$ and the vertices of $T$ correspond to $V_{1}$, it follows that $\psi(u)=u$ for every $u\in V_{1}$. By Lemma 2.1, we conclude that $\psi=id_{C(G)}$.
@@
  1. Case A: Suppose $G$ has a central vertex $v_{0}$. If $S_{2}(v_{0})\neq\emptyset$, we consider $G$ rooted in $v_{0}$ and apply the coloring argument of Case (ii), which fixes vertices successively according to their distance from $v_{0}$. If $S_{2}(v_{0})=\emptyset$, then $G\cong K_{1,n-1}$. In this case, it is known that
    $D^{\prime}(C(K_{1,n-1}))=\left\lceil\sqrt{n-1}\right\rceil=\left\lceil\sqrt{\Delta(K_{1,n-1})}\right\rceil.$
    Thus, the desired bound holds.
  2. Case B: If $G$ has a central edge $e_{0}=xy$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$. If we put different colors on the edges $xw_{xy}$ and $w_{xy}y$, then $x$ and $y$ are fixed by every automorphism. We color $T_{i}$ using the same method as in Case A for $i\in\{1,2\}$.
Proposition 3.6.
For all integers $m,n\geq 3$,
$D^{\prime}(C(K_{m,n}))\leq\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))+1.$
Proof.
Let $V=\{v_{1},\dots,v_{n}\}$ and $U=\{u_{1},\dots,u_{m}\}$ be the bipartitions of $K_{m,n}$, and let $G=C(K_{m,n})$. Then $G[V]\cong K_{n}$ and $G[U]\cong K_{m}$ where $G[V]$ and $G[U]$ are the induced subgraphs of $G$ with respect to $V$ and $U$ respectively. Let $G_{n}:=G[V]$ and $G_{m}:=G[U]$ (see Figure 4).
[…]
Figure 4. A schematic representation of $C(K_{n,m})$. Representative edges of the cliques $K_{n}$ and $K_{m}$ are drawn, and each edge $v_{i}u_{j}$ is subdivided by a vertex $w_{v_{i}u_{j}}$.
Suppose that $f_{n}:E(G_{n})\to\{1,\dots,D^{\prime}(K_{n})\}$ and $f_{m}:E(G_{m})\to\{1,\dots,D^{\prime}(K_{m})\}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$, respectively. Let $d=\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))$. Define a coloring $f:E(G)\to\{1,\dots,d+1\}$ as follows:
[…]
Claim 3.7.
$f$ is a distinguishing edge coloring of $G$.
Proof.
Let $\phi\in\mathrm{Aut}(G)$ be an automorphism that preserves $f$. By Lemma 2.2, $\phi(V\cup U)=V\cup U$ and $\phi(V_{2})=V_{2}$. For each $v_{i}\in V$, all edges of the form $v_{i}w_{v_{i}u_{j}}$ are colored $d+1$ with respect to $f$, whereas for each $u_{j}\in U$, no edge of the form $w_{v_{i}u_{j}}u_{j}$ is colored $d+1$. Thus, the property
[…]
characterizes vertices of $V$. Since $\phi$ preserves $f$, it follows that $\phi(V)=V$ and $\phi(U)=U$. Define $\psi_{V}:=\phi\restriction V$ and $\psi_{U}:=\phi\restriction U$, i.e., the restrictions of $\phi$ to $V$ and $U$ respectively. It is clear that $\psi_{V}\in Aut(G_{n})$ and $\psi_{U}\in Aut(G_{m})$. [^2] Let $v_{i}v_{j}\in E(G_{n})$. Then $v_{i}v_{j}\in E(G)$ and both endpoints $v_{i}$ and $v_{j}$ lie in $V$. Since $\phi\in Aut(G)$, we have $\phi(v_{i})\phi(v_{j})\in E(G)$. Since $\phi(V)=V$, it follows that $\phi(v_{i}),\phi(v_{j})\in V$, and hence $\phi(v_{i})\phi(v_{j})\in E(G_{n})$. Thus, $\psi_{V}$ preserves adjacency in $G_{n}$, and hence $\psi_{V}\in Aut(G_{n})$. Similarly, $\psi_{U}\in Aut(G_{m})$.
Fix any $v_{i}v_{j}\in E(G_{n})$. Then $\psi_{V}(v_{i})\psi_{V}(v_{j})\in E(G_{n})$ where $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$ belong to $V$. Since $\psi_{V}\in Aut(G_{n})$ and automorphisms preserve adjacency, the endpoints of $\psi_{V}(v_{i}v_{j})$ are $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$. Now $G_{n}\cong K_{n}$ is a simple graph. Thus, we have
$\psi_{V}(v_{i}v_{j})=\psi_{V}(v_{i})\psi_{V}(v_{j}).$
Since $\phi$ preserves $f$, we obtain
$f_{n}(\psi_{V}(v_{i}v_{j}))=f_{n}(\psi_{V}(v_{i})\psi_{V}(v_{j}))=f(\psi_{V}(v_{i}v_{j}))=f(\phi(v_{i}v_{j}))=f(v_{i}v_{j})=f_{n}(v_{i}v_{j}).$
Thus, $\psi_{V}$ preserves $f_{n}$. Similarly, $\psi_{U}$ preserves $f_{m}$. We know $f_{n}$ and $f_{m}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$ respectively. Thus, $\psi_{V}$ and $\psi_{U}$ fixes all vertices of $V$ and $U$ respectively. Consequently, $\phi$ fixes $V\cup U$ pointwise. By Lemma 2.1, we conclude that $\phi=id_{C(K_{m,n})}$. ∎
@@
In view of Fact 3.1(1), it suffices to show that $D(C(G))\leq D^{\prime\prime}(G)$. Let $f:V(G)\cup E(G)\to\{1,...,D^{\prime\prime}(G)\}$ be a total coloring of $G$ (not necessarily proper) that is preserved only by $id_{G}$. We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{xy})=f(xy)$ for all $w_{xy}\in V_{2}$. We claim that $\tilde{f}$ is preserved only by $id_{C(G)}$. Let $\psi$ be a $\tilde{f}-$preserving automorphism of $C(G)$. If $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$ where $\pi_{\alpha}$ is the induced automorphism of $\alpha\in Aut(G)$, which is defined in the proof of Lemma 2.3, then $H=Aut(C(G))\cong Aut(G)$ (see the proof of Lemma 2.3). Thus, $\psi=\pi_{\alpha}$ for some $\alpha\in Aut(G)$ where $\psi(V_{1})=V_{1}$ and $\psi(V_{2})=V_{2}$. Since $\psi$ is a $\tilde{f}-$preserving automorphism of $C(G)$, we can see that $\alpha$ preserves $f$ for the following reasons:
@@
    If $v\in V_{1}$, then
    $\tilde{f}(\psi(v))$ $=\tilde{f}(v)$
    $\implies f(\psi(v))=f(v)$
    $\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v).$
@@
    If $w_{xy}\in V_{2}$, then
    $\tilde{f}(\psi(w_{xy}))$ $=\tilde{f}(w_{xy})$
    $\implies\tilde{f}(\pi_{\alpha}(w_{xy}))=f(xy)$
    $\implies\tilde{f}(w_{\alpha(x)\alpha(y)})=f(xy)$
    $\implies f(\alpha(x)\alpha(y))=f(xy)$
    $\implies f(\alpha(xy))=f(xy).$
@@
For any integer $n\geq 1$, we have
$D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil.$
Thus, there exist graphs for which the bounds mentioned in Theorems 3.4 and 3.8 cannot be improved (see Fig. 5).
@@
Let $E$ be the set of all endline edges of $G^{+}$.
(1). By Facts 3.1(4) and 2.5 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
Case(i): $G\not\cong C_{n}$ for any $n\geq 3$. We show that $D^{\prime}(G^{+})\leq D^{\prime}(G)$. Then $D^{\prime}(G^{+})\leq\Delta(G)$ by Fact 3.1(5). Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$. Define the mapping $g^{\prime}:E(G^{+})\rightarrow\{1,...,D^{\prime}(G)\}$ as follows:
@@
Case(ii): $G\cong C_{n}$ for some $n\geq 3$.
Claim 3.11.
$D^{\prime}(G^{+})=2$.